Need help with the following:

Determine whether the graph of y = |x| + 1 is symmetric with respect to the x-axis, the y-axis, or the origin (or none).

Answer by venugopalramana(3286)

Determine whether the graph of y=F(X) = |x| + 1
F(+X)=X+1
F(-X)=|-X|+1=X+1...HENCE
F(X)=F(-X)
SO THE GRAPH IS SYMMETRIC W.R.T.X=0 OR Y AXIS.
is symmetric with respect to the x-axis, the y-axis, or the origin (or none).